OpenStudy (anonymous):

Need help one two questions... will reward with fan & medal and its kinda urgent. Thanks its really appreciated! 1. Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). 2. Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph.

OpenStudy (anonymous):

Here is the graph for #2 file:///Users/sydney/Desktop/03_07_18.jpg

OpenStudy (triciaal):

change to form y = m x + b m is the slope parallel lines have the same slope for 2 points on the line (x, y) and (x1, y1) m = (y1 - y)/ (x1 - x) use (-3, 4) as (x1, y1)

OpenStudy (triciaal):

2) parallel lines have the same slope, m point on the line (x, y) and given (3, -1) y + 1 = m( x -3) y = mx -3*m -1 use the m from above

OpenStudy (anonymous):

Thanks but I am still confused on the first one?

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

I didn't realize you have a graph for #2 m is the slope of the line in the given graph that I didn't know about slope = change in y/change in x

OpenStudy (anonymous):

Oh I was confused for a second. So, would you solve with that formula? @triciaal

OpenStudy (anonymous):

so what is the answer for #1? I'm also confused for what you put sorry haha

OpenStudy (triciaal):